(1 point) A typical sheet of weighing paper is 3.8 inches by 3.6 inches and weighs 0.45 g. What is its absorber thickness in g/cm2?
Answer #1

given that width s=3.8 inches=9.652cm

                   length l = 3.6 inches=9.144 cm

hence area is A=l*s

                         =88.258 cm^2

mass m=0.45g

   absorber thickness =0.45/88.258

                              =5.0987*10^-3 g/cm^2